A metal cylinder of mass 4kg vibrates harmonically with an amplitude of 0.25m and period of 2second. Calculate 1)Frequency 2)its maximum speed and speed when y=0.15m


1
Expert's answer
2021-01-24T14:23:50-0500

Answer

1) frequency is inverse of time period

"f=\frac{1}{T}=\frac{1}{2}=0.5Hz."

2) SHM equation is given by

"y=y_0sin\omega t"

Velocity

"v=v_0cos\omega t"

Where "v_0" is "y_0 \omega"

Maximum velocity

"v=v_0=y_0\omega"

"=0.15\times2\pi\times0.5"

=0.471 m/s